# Anytime-valid t-tests and confidence sequences for Gaussian means with unknown variance

### v4

@@ 2.4 Sequential t-Test and t-Confidence Sequences
Different objectives are pursued for the sequential Gaussian t-test problem. Early authors have studied the problem of testing the null $\mathcal{N}_{\mu=0}$ (or more generally $\mathcal{N}_{\mu/\sigma=\theta_{0}}$) against the alternative $\mathcal{N}_{\mu/\sigma=\theta_{1}}$ [Rushton, 1950, Ghosh, 1960, Sacks, 1965], and its scale-invariant nature translates into the group-invariant setting as the problem of testing orbital nulls and alternatives [Pérez-Ortiz et al., 2022].
In this paper, we shall roughly follow the same routine as Lai [1976] and be less concerned with the alternatives, chiefly focusing instead on the construction of (1) test processes for the null $\mathcal{N}_{\mu=0}$ (and study their behavior under both null and non-null underlying distributions); (2) Confidence sequences for the population mean $\mu$ over the set of all Gaussian distributions $\mathcal{N}$, which we shall often shorthand as “t-confidence sequences”, or “t-CSs” for brevity. One may note that (1) is sufficient to produce tests for arbitrary point location null $\mathcal{N}_{\mu=\mu_{0}}$ by shifting, and consequently (2) by Lemmas 2.1 and 2.2 and inversion. Occasionally we are interested in test processes for the one-sided null $\mathcal{N}_{\mu\leqslant 0}$ as these give rise to one-sided tests and CSs. In fact, we shall later construct test processes for $\mathcal{N}_{\mu/\sigma=\theta_{0}}$ which is the strongest statistical quantity in the setting, as it leads to both CSs and tests for the mean $\mu$ and the normalized mean $\mu/\sigma$.
@@ 2.5 Likelihood Ratio Martingales
We shall use two stronger forms of Lemma 2.3, stated below and proved in Section C.1. First, the distribution on the numerator, $Q$, can be varying and depend on previous observations.
@@
Let $P$ be a probability distribution on $\mathbb{R}$. For each $n\geqslant 1$ let $Q_{n}:\Omega\times\mathcal{B}(\mathbb{R})\to\mathbb{R}$ be a Markov kernel such that: 1) $Q_{n}(\,\cdot\,,B)$ is $\mathcal{F}_{n-1}$-measurable for all $B\in\mathcal{B}(\mathbb{R})$, and 2) $Q_{n}(\omega,\,\cdot\,)\ll P$ for all $\omega\in\Omega$. Then, the process
@@
We shall demonstrate how the method of universal inference [Wasserman et al., 2020] leads to e-processes and confidence sequences for the t-test. Throughout this section, for $n\geqslant 0$, we denote by $\widetilde{\mu}_{n}$ and $\widetilde{\sigma}_{n}$ any point estimators for $\mu$ and $\sigma$ adapted to the canonical filtration $\mathcal{F}_{n}$ (i.e. based on $X_{1},\dots,X_{n}$). For example, they can simply be the sample mean $\widehat{\mu}_{n}$ and sample standard deviation $\sqrt{v_{n}}$; we may also use the Bayesian approach by sequentially updating them as the posterior mean from, say, a normal-inverse-gamma prior on $(\mu,\sigma)$111which is the conjugate prior of normal distributions when both of $\mu$ and $\sigma$ are unknown. The method of sequential universal inference [Wasserman et al., 2020, Section 8] implies the following plug-in likelihood ratio martingale for a fixed Gaussian distribution:
@@
Under any $P=N_{\mu,\sigma^{2}}$, suppose there is a $\gamma>0$ such that $\{\widetilde{\mu}_{n}\}$ converges to $\mu$ in $L^{3}$ with rate $\mathrm{E}_{P}(|\widetilde{\mu}_{n}-\mu|^{3})\lesssim n^{-\gamma}$, $\{\widetilde{\sigma}_{n}^{-2}\}$ converges to $\sigma^{-2}$ both in $L^{2}$ with rate $\mathrm{E}_{P}(\widetilde{\sigma}_{n}^{-2}-\sigma^{-2})^{2}\lesssim n^{-\gamma}$ and almost surely, and has uniformly bounded 3rd moment $\limsup_{n}\mathrm{E}_{P}(\widetilde{\sigma}_{n}^{-6})<\infty$. Then,
@@
The assumptions on the point estimators are mild, as they are satisfied by the empirical mean and variance due to the moment properties of the inverse-$\chi^{2}$ distribution, as well as their smoothed or Bayesian extensions (posterior means under reasonable priors). The exponential growth of $\{R_{n}\}$ under the alternative $\mathcal{N}_{\mu\neq 0}$ is in contrast to its restrained behavior under the null $\mathcal{N}_{\mu=0}$, characterized by Ville’s inequality for e-processes (Lemma 2.1).
@@
Under any $N_{\mu,\sigma^{2}}$, suppose the point estimators $\{\widetilde{\mu}_{n}\}$ and $\{\widetilde{\sigma}_{n}^{-2}\}$ satisfy the same assumptions as Proposition 3.3. Then,
@@ 4.1 Lai’s [1976] Confidence Sequence
Choose a starting time $m\geqslant 2$ and a constant $a>0$. Recall that $v_{n}=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\widehat{\mu}_{n})^{2}$. Further, define
@@ 4.2 Scale Invariant Filtration
The filtration $\{\mathcal{F}_{n}^{*}\}$ contains all the information up to time $n$ about the relative sizes of the observations. For example, $\{\max\{X_{1},\dots,X_{4}\}\geqslant 2X_{3}\}\in\mathcal{F}_{4}$ while $\{\max\{X_{1},\dots,X_{4}\}\geqslant 2\}\notin\mathcal{F}_{4}$. The reader may recall the exchangeable sub-$\sigma$-algebra in the theory of exchangeability and backwards martingales such as in Klenke [2013, Chapter 12] for an analogy. Actually, when each of the observations $X_{1},X_{2},\dots$ is non-zero, it has the following clean expression.
@@
In our scenario of application, the t-test under Gaussian distribution, data are not non-zero but almost surely non-zero. We do not distinguish (25) and (26) as the definition of the scale invariant filtration because processes adapted to it are often derived by manipulations of $X_{1}/|X_{1}|,\dots,X_{n}/|X_{1}|$, which we call the scale invariant reduction of the observations, ignoring the $N_{\mu,\sigma^{2}}$-negligible event $\{X_{1}=0\}$. Indeed, Lai [1976] defines such filtration via (26) for the t-test case. However, we remark that one would need the more general (25) in Definition 4.2 for the case when $X_{1}\neq 0$ does not hold almost surely (e.g. when exploring a nonparametric extension of our set-up and methods).
As a counter-example, one may verify that the e-processes for t-test that we have previously introduced via universal inference, $\{R_{n}\}$ in (10) and $\{R_{n}^{-}\}$ in (15), are not adapted to $\{\mathcal{F}_{n}^{*}\}$. They would have been if the plug-in estimates $\widetilde{\mu}_{i-1}$ and $\widetilde{\sigma}_{i-1}$ were allowed to be homogeneous of degree 1. However, as $\widetilde{\mu}_{0}$ and $\widetilde{\sigma}_{0}$ must be constants, this cannot be the case. In fact, when $n=1$, these two e-processes are necessarily non-constant, and thus are not $\mathcal{F}_{1}^{*}$-measurable.
4.3 Scale Invariant Likelihood Ratios
@@
Let $V_{n}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the function
$V_{n}(x_{1},\dots,x_{n})=(x_{1}/|x_{1}|,x_{2}/|x_{1}|,\dots,x_{n}/|x_{1}|).$ (28)
(It does not matter how $V_{n}$ is defined when $x_{1}=0$.)
Let $Q_{\mu,\sigma^{2}}^{n}$ be the push-forward measure of $P_{\mu,\sigma^{2}}^{\otimes n}$ under the map $V_{n}$. Let $\eta^{\langle n-1\rangle}$ be the measure on {±$\{\pm 1\}\times\mathbb{R}^{n-1}$ that charges both $\{1\}\times\mathbb{R}^{n-1}$ and $\{-1\}\times\mathbb{R}^{n-1}$ with $\eta^{\otimes(n-1)}$. Then,
@@
Where $\mathrm{J}$ is the Jeffreys prior over $\mathcal{P}$ with density $\frac{1}{\sigma(P)}$. This improper prior is known in the Bayesian literature as an uninformative prior on the scale parameter. We have shown above that taking the mixture (in some sense) with the Jeffreys prior $\int_{\tau>0}(\dots)\frac{\mathrm{d}\tau}{\tau}$ of likelihood ratios is equivalent to taking the likelihood ratio of the scale invariant reduction $X_{2}/|X_{1}|,\dots,X_{n}/|X_{1}|$. We shall make further remarks on how our approaches are related to previous Bayesian work with Jeffreys prior in Section 4.6.
@@ 4.4 An Extended Test Martingale for t-Test
Under any $N_{\mu,\sigma^{2}}$,
@@
We remark that the reduction of filtration from $\{\mathcal{F}_{n}\}$ to $\{\mathcal{F}^{*}_{n}\}$ does limit the stopping rules that are safe when working with the test processes. Stopping on a scale variant event would violate the safety of $\{H_{n}\}$. However, the effect of filtration reduction seems to disappear after applying extended Ville’s inequality to obtain a CS. This is because Ville’s inequality or extended Ville’s inequality, at a fundamental level, makes a claim only about the first exit time from $[0,\varepsilon]$ of the process (NSM or ENSM) — a stopping time on the process itself, one adapted to the canonical filtration of the process which is even coarser than $\{\mathcal{F}^{*}_{n}\}$.
@@ 4.5 Classical Test Martingales for t-Test
While the test martingale (36) works without any limitation on $n$ or $c$, the CS (38) is non-trivial only when the $\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}(n+c^{2})-c^{2}$ term in the denominator is positive. If a CI at time $n$, or the width of the CS at some fixed time $n$ is of interest, one can take $c^{2}$ such that
@@
which is always finite and has a growth rate of $\alpha^{-1/(n-1)}$ as $\alpha\to 0$ while keeping the sample fixed, a slightly worse rate compared to the $\alpha^{-1/n}$ of the universal inference CS, and at least as good as the $\alpha^{-m/n(m-1)}$ of Lai’s original CS since $m\leqslant n$. On the other hand, if $c$ and $n$ are both fixed, the radius grows like $(\alpha-\alpha_{\star}(n))^{-1/n}$ as $\alpha\downarrow\alpha_{\star}(n)=\exp({\Theta}(-n\log n))$.
For a fixed $\alpha$ and a fixed $c$, the CS is non-trivial on $n\geqslant n_{0}$ for some $n_{0}$, instead of all $n\geqslant 1$. For example, when $c=0.01$ and $\alpha=0.05$, the range of (38) is $n\geqslant 3$. Thus, the starting time $m$ in Theorem 4.1 seems to be avoided by switching from the extended Ville’s inequality to the classical Ville’s inequality, but it remains in another form. In this case, if $n\to\infty$, $\operatorname{radius}_{n}$ shrinks as a function of $n$ at the rate of
@@
Under any $N_{\mu,\sigma^{2}}$,
@@
The NM $\{G_{n}^{(c)}\}$ thus have the same asymptotic properties as the ENSM $\{H_{n}\}$, but the free parameter $c$, which is absence for the ENSM, does introduce a difference that emerges only non-asymptotically. To wit, if one fixes $n$ and the data-dependent quantity $T^{2}_{n-1}$, the value of $G_{n}^{(c)}$ would approach 0 (hence the power vanishes) if $c$ is too large or too small.
@@
We also note that Lindon et al. [2022, Proposition 3.7] concurrently derive the same t-test martingale and confidence sequence as our Theorem 4.9, a special case of their F-test martingale. Our paper, however, is concerned much more with comparing the Gaussian mixture with Lai’s flat mixture in theory and practice, with understanding the tradeoffs between test martingales in the reduced filtration and e-processes in the original filtration for this fundamental problem, as well as deriving their (often suboptimal) dependence on $\alpha$.
In place of the full Gaussian prior, a half-Gaussian prior mixture will give us a “semi-one-sided” test process in the sense that it is valid only for the point null $\mathcal{N}_{\mu=0}$ (unlike Theorem 3.4) but still leads to a one-sided confidence sequence.
@@
$G_{n}^{(c-)}=2\sqrt{\frac{c^{2}}{n+c^{2}}}\left(\left(1-\frac{S_{n}^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}-\left(1-\frac{(S_{n}\wedge 0)^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}\right)=2G_{n}^{(c)}-2G_{n}^{(c)}|_{(S_{n}\leftarrow S_{n}\wedge 0)}.$ (44)

### v5

@@ 2.5 Sequential t-Test and t-Confidence Sequences
Different objectives are pursued for the sequential Gaussian t-test problem. Early authors have studied the problem of testing the null $\mathcal{N}_{\mu=0}$ (or more generally $\mathcal{N}_{\mu/\sigma=\theta_{0}}$) against the alternative $\mathcal{N}_{\mu/\sigma=\theta_{1}}$ [Rushton, 1950, Ghosh, 1960, Sacks, 1965], and its scale-invariant nature translates into the group-invariant setting as the problem of testing orbital nulls and alternatives [Pérez-Ortiz et al., 2024].
In this paper, we shall roughly follow the same routine as Lai [1976] and aim at power against a broader range of alternatives, chiefly focusing instead on the construction of (1) test processes for the null $\mathcal{N}_{\mu=0}$ (and study their behavior under both null and non-null underlying distributions); (2) Confidence sequences for the population mean $\mu$ over the set of all Gaussian distributions $\mathcal{N}$, which we shall often shorthand as “t-confidence sequences”, or “t-CSs” for brevity. One may note that (1) is sufficient to produce tests for arbitrary point location null $\mathcal{N}_{\mu=\mu_{0}}$ by shifting, and consequently (2) by Lemmas 2.1 and 2.2 and inversion. Occasionally we are interested in test processes for the one-sided null $\mathcal{N}_{\mu\leqslant 0}$ as these give rise to one-sided tests and CSs. In fact, we shall later construct test processes for $\mathcal{N}_{\mu/\sigma=\theta_{0}}$ which is the strongest statistical quantity in the setting, as it leads to both CSs and tests for the mean $\mu$ and the normalized mean $\mu/\sigma$.
@@ 2.6 Likelihood Ratio Martingales
Recalling Definition 2.3, the e-power of $\{L_{n}\}$ equals the Kullback-Leibler divergence $D_{\operatorname{KL}}(Q\|P)$. The growth of $L_{n}=\exp(nD_{\operatorname{KL}}(Q\|P)+o_{Q}(n))$ is optimal in this point versus point case, matching the ${\Theta}\left(\log(1/\alpha)D_{\operatorname{KL}}(Q\|P)^{-1}\right)$ lower bound in the expected rejection time with type 1 error rate $\alpha$. See e.g. Garivier et al. [2019, Section 2] and Kaufmann et al. [2016, Lemma 1] for this lower bound in the equivalent bandit formulation.
We shall use two stronger forms of Lemma 2.4 stated below and proved in Section C.1. First, the distribution on the numerator, $Q$, can be varying and depend on previous observations.
@@
Let $P$ be a probability distribution on $\mathbb{R}$. For each $n\geqslant 1$ let $Q_{n}:\Omega\times\mathcal{B}(\mathbb{R})\to\mathbb{R}$ be a Markov kernel such that: 1) $Q_{n}(\,\cdot\,,B)$ is $\mathcal{F}_{n-1}$-measurable for all $B\in\mathcal{B}(\mathbb{R})$, and 2) $Q_{n}(\omega,\,\cdot\,)\ll P$ for $P$-almost all $\omega\in\Omega$. Then, the process
@@
We shall demonstrate how the method of universal inference [Wasserman et al., 2020] leads to e-processes and confidence sequences for the t-test. Throughout this section, for $n\geqslant 0$, we denote by $\widetilde{\mu}_{n}$ and $\widetilde{\sigma}_{n}$ any point estimators for $\mu$ and $\sigma$ adapted to the canonical filtration $\mathcal{F}_{n}$ (i.e. based on $X_{1},\dots,X_{n}$). For example, they can simply be the sample mean $\widehat{\mu}_{n}$ and sample standard deviation $s_{n}$; we may also use the Bayesian approach by sequentially updating them as the posterior mean or maximum a posteriori from, say, a normal-inverse-gamma prior [^3] The normal-inverse-gamma distribution serves as the conjugate prior for the Gaussian family when both of $\mu$ and $\sigma$ are unknown. See Section 5.2 for more description. on $(\mu,\sigma^{2})$. The method of sequential universal inference [Wasserman et al., 2020, Section 8] implies the following plug-in likelihood ratio martingale for a fixed Gaussian distribution:
@@
Under any distribution $P$ with mean $\mu$ and variance $\sigma^{2}$, for example $N_{\mu,\sigma^{2}}$, suppose there is a $\gamma>0$ such that $\{\widetilde{\mu}_{n}\}$ converges to $\mu$ in $L^{3}$ with rate $\mathrm{E}_{P}(|\widetilde{\mu}_{n}-\mu|^{3})\lesssim n^{-\gamma}$, $\{\widetilde{\sigma}_{n}^{-2}\}$ converges to $\sigma^{-2}$ both in $L^{2}$ with rate $\mathrm{E}_{P}(\widetilde{\sigma}_{n}^{-2}-\sigma^{-2})^{2}\lesssim n^{-\gamma}$ and almost surely, and has uniformly bounded 3rd moment $\limsup_{n}\mathrm{E}_{P}(\widetilde{\sigma}_{n}^{-6})<\infty$. Then,
@@
Note that we place no Gaussian assumption on the true alternative distribution $P$ for this e-power statement, only that the plug-in point estimators need to satisfy some convergence assumptions. These assumptions, we note, are mild, as they are satisfied by the empirical mean and variance under $N_{\mu,\sigma^{2}}$ due to the moment properties of the inverse-$\chi^{2}$ distribution, as well as their smoothed or Bayesian extensions (posterior means or maximum a posteriori under reasonable priors). The exponential growth of $\{R_{n}\}$ under the alternative $\mathcal{N}_{\mu\neq 0}$ is in contrast to its restrained behavior under the null $\mathcal{N}_{\mu=0}$, characterized by Ville’s inequality for e-processes (Lemma 2.1).
@@
Under any distribution $P$ with mean $\mu$ and variance $\sigma^{2}$, for example $N_{\mu,\sigma^{2}}$, suppose the point estimators $\{\widetilde{\mu}_{n}\}$ and $\{\widetilde{\sigma}_{n}^{-2}\}$ satisfy the same assumptions as Proposition 3.3. Then,
@@ 4.1 Lai’s [1976] Confidence Sequence
Choose a starting time $m\geqslant 2$ and a constant $a>0$. Recall that $s_{n}^{2}=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\widehat{\mu}_{n})^{2}$. Further, define
@@ 4.2 Scale Invariant Filtration
The filtration $\{\mathcal{F}_{n}^{*}\}$ contains all the information up to time $n$ about the relative sizes of the observations. For example, $\{\max\{X_{1},\dots,X_{4}\}\geqslant 2X_{3}\}\in\mathcal{F}_{4}$ while $\{\max\{X_{1},\dots,X_{4}\}\geqslant 2\}\notin\mathcal{F}_{4}$. The reader may recall the exchangeable sub-$\sigma$-algebra in the theory of exchangeability and backwards martingales such as in Klenke [2013, Chapter 12] for an analogy. The reader might also find the following analogy helpful in understanding the intuition: Imagine in prehistoric times when units of length were yet to be invented, our ancestors collected a sequence of leaves and fathomed their sizes. Due to the lack of units however, no absolute real-valued sizes (in inches, say) but only ratios of sizes relative to each other (or to the first leaf) were recorded. Actually, when $X_{1}$ is non-zero, $\mathcal{F}_{n}^{*}$ has the following clean expression.
@@
In our scenario of application, the t-test under Gaussian distribution, data are not non-zero but almost surely non-zero. We do not distinguish (24) and (25) as the definition of the scale-invariant filtration because processes adapted to it are often derived by manipulations of $X_{1}/|X_{1}|,\dots,X_{n}/|X_{1}|$, which we call the scale invariant reduction of the observations, ignoring the $N_{\mu,\sigma^{2}}$-negligible event $\{X_{1}=0\}$. Indeed, Lai [1976] defines such filtration via (25) for the t-test case. However, we remark that one would need the more general (24) in Definition 4.2 for the case when $X_{1}\neq 0$ does not hold almost surely (e.g. when exploring a nonparametric extension of our set-up and methods).
As a counter-example, one may verify that the e-processes for t-test that we have previously introduced via universal inference, $\{R_{n}\}$ in (9) and $\{R_{n}^{-}\}$ in (14), are not adapted to $\{\mathcal{F}_{n}^{*}\}$. They would have been if the plug-in estimates $\widetilde{\mu}_{i-1}$ and $\widetilde{\sigma}_{i-1}$ were allowed to be homogeneous of degree 1. However, as $\widetilde{\mu}_{0}$ and $\widetilde{\sigma}_{0}$ must be constants, this cannot be the case. In fact, when $n=1$, these two e-processes are necessarily non-constant, and thus are not $\mathcal{F}_{1}^{*}$-measurable. In what follows, we shall present numerous test processes on the scale invariant filtration $\{\mathcal{F}^{*}_{n}\}$, beginning with some generality. A close-up discussion on this intriguing issue follows afterward, in Section 4.6.5.
4.3 Scale Invariant Likelihood Ratios for Location-Scale Families
@@
Let $\varphi_{n}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the function
$\varphi_{n}(x_{1},\dots,x_{n})=(x_{1}/|x_{1}|,x_{2}/|x_{1}|,\dots,x_{n}/|x_{1}|).$ (27)
(It does not matter how $\varphi_{n}$ is defined when $x_{1}=0$.)
Let $Q_{\mu,\sigma^{2}}^{n}$ be the push-forward measure of $P_{\mu,\sigma^{2}}^{\otimes n}$ under the map $\varphi_{n}$. Let $\eta^{\langle n-1\rangle}$ be the measure on {±$\{\pm 1\}\times\mathbb{R}^{n-1}$ that charges both $\{1\}\times\mathbb{R}^{n-1}$ and $\{-1\}\times\mathbb{R}^{n-1}$ with $\eta^{\otimes(n-1)}$. Then,
@@
Where $\mathrm{J}$ is the Jeffreys prior over $\mathcal{P}$ with density $\frac{1}{\sigma(P)}$. This improper prior is known in the Bayesian literature as an uninformative prior on the scale parameter. We have shown above that taking the mixture (in some sense) with the Jeffreys prior $\int_{\tau>0}(\dots)\frac{\mathrm{d}\tau}{\tau}$ of likelihood ratios is equivalent to taking the likelihood ratio of the scale invariant reduction $X_{2}/|X_{1}|,\dots,X_{n}/|X_{1}|$. Indeed, our treatment for the location-scale family $\{P_{\mu,\sigma^{2}}\}$ so far in this subsection to compute the density of $Q^{n}_{\mu,\sigma^{2}}$ is a special case of group invariant models formulated by Pérez-Ortiz et al. [2024, Section 2.2], where the Jeffreys prior $J(\mathrm{d}P)$ with density $\sigma^{-1}(P)$ arises as the Haar measure of the group $(\mathbb{R}^{+},\times)$, and the function $V_{n}$ corresponds to the “maximally invariant statistic” under the group action. The exact relation of our work to that of Pérez-Ortiz et al. [2024] shall be discussed in detail in Section 4.6.3. We shall make further remarks on how our approaches are related to previous Bayesian work with Jeffreys prior in Section 4.6.4.
@@ 4.4 An Extended Test Martingale for t-Test
Under any distribution with mean $\mu$ and variance $\sigma^{2}$, for example $N_{\mu,\sigma^{2}}$,
@@ 4.5 Classical Test Martingales for t-Test
The tuning parameter $c^{2}$ of prior precision, we note, does not affect the safety (the sequential test having type 1 error rate within $\alpha$) or coverage (the CS covering $\mu$ with at least $1-\alpha$ probability) of our method, but does affect the power of the sequential test and the tightness of the confidence sequence. First, considering the growth of the e-process under the alternative, since the mixing distribution $N_{0,c^{-2}}$ is put over all possibilities of the true effect size $\theta=\mu/\sigma$, it would be ideal if $N_{0,c^{-2}}$ “covers” the true underlying effect size with adequate weight. Therefore, if one has acquired prior knowledge of the rough scale of the true effect size $\theta$, matching $c^{-1}$ to the same scale can improve power, as we shall demonstrate numerically in Section 5.2. However, as we shall see soon in Proposition 4.10, in the large sample regime different choices of $c^{2}$ lead to the same e-power under a fixed alternative distribution. On the other hand, the tuning parameter $c^{2}$ plays a distribution-free role on the tightness of the $(1-\alpha)$-CS, as the expression for the radius (36) contains decoupled $(c,\alpha,n)$-dependent and data-dependent terms. In this case, difference choices of $c^{2}$ optimize the radius of the CS at different times $n$. As discussed in a similar situation of the Gaussian mixture Z-test CS by Howard et al. [2021, Section 3.5], there is a trade-off in “making a bound tighter for some range of times requires making it looser at other times.” We shall numerically demonstrate this in our case in Section 5.4. More specifically, to approximately minimize $\operatorname{radius}_{n}$ at some fixed $n$, one can take $c^{2}$ such that
@@
which is always finite and has a growth rate of $\alpha^{-1/(n-1)}$ as $\alpha\to 0$ while keeping the sample fixed, a slightly worse rate compared to the $\alpha^{-1/n}$ of the universal inference CS, and at least as good as the $\alpha^{-m/n(m-1)}$ rate of Lai’s original CS since $m\leqslant n$. On the other hand, if $c$ and $n$ are both fixed, the radius grows like $(\alpha-\alpha_{\star}(n))^{-1/n}$ as $\alpha\downarrow\alpha_{\star}(n)=\exp({\Theta}(-n\log n))$.
While the test martingale (35) works without any limitation on $n$ or $c$, the CS (37) is non-trivial only when the $\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}(n+c^{2})-c^{2}$ term in the denominator is positive. For a fixed $\alpha$ and a fixed $c$, the CS is non-trivial on $n\geqslant n_{0}$ for some $n_{0}$, instead of all $n\geqslant 1$. For example, when $c=0.01$ and $\alpha=0.05$, the range of (37) is $n\geqslant 3$. Thus, the starting time $m$ in Theorem 4.1 seems to be avoided by switching from the extended Ville’s inequality to the classical Ville’s inequality, but it remains in another form. In this case, if $n\to\infty$, $\operatorname{radius}_{n}$ shrinks as a function of $n$ at the rate of
@@
Under any distribution with mean $\mu$ and variance $\sigma^{2}$, for example $N_{\mu,\sigma^{2}}$,
@@
The NM $\{G_{n}^{(c)}\}$ thus have the same asymptotic properties as the ENSM $\{H_{n}\}$, but the free parameter $c$, which is absent for the ENSM, does introduce a difference that emerges only non-asymptotically. To wit, if one fixes $n$ and the data-dependent quantity $T^{2}_{n-1}$, the value of $G_{n}^{(c)}$ would approach 0 (hence the power vanishes) if $c$ is too large or too small.
@@
We also note that Lindon et al. [2022, Proposition 3.7] concurrently derive the same t-test martingale and confidence sequence as our Theorem 4.9, a special case of their F-test martingale. Our paper, however, is concerned much more with comparing the Gaussian mixture with Lai’s flat mixture in theory and practice, with understanding the tradeoffs between test martingales in the reduced filtration and e-processes in the original filtration for this fundamental problem, as well as deriving their dependence on $\alpha$.
In place of the full Gaussian prior, a half-Gaussian prior mixture over $\mu/\sigma>0$ will give us a “semi-one-sided” test process in the sense that it is valid only for the point null $\mathcal{N}_{\mu=0}$ (unlike Theorem 3.4) but still leads to a test powerful against $\mathcal{N}_{\mu>0}$ and a one-sided confidence sequence.
@@
$\scriptsize G_{n}^{(c-)}=2\sqrt{\frac{c^{2}}{n+c^{2}}}\left(\left(1-\frac{S_{n}^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}-\left(1-\frac{(S_{n}\wedge 0)^{2}}{(n+c^{2})V_{n}}\right)^{-n/2}\right)=2G_{n}^{(c)}-2G_{n}^{(c)}|_{(S_{n}\leftarrow S_{n}\wedge 0)}.$ (43)